# Towards Open Vocabulary Learning: A Survey

### v1

@@ 4 Benchmark Results
In this section, we systematically compare the different settings and methods in each task. First, we introduce the datasets used in the evaluation. Next, we present the details of each setting, separately. Then, we compare the results in detailed tables. We list the top performance models for reference.
@@
TABLE III: Open vocabulary object detection results on the COCO dataset. The evaluation metric is box mAP at IoU threshold 0.5 for novel and base classes, respectively.
@@
Method Vision Backbone Text Model Trainable Backbone Detector Extra Data $AP_{novel}$ $AP_{base}$ $AP_{all}$ OVR-CNN [12] (CVPR ’21) R50-C4 BERT Faster R-CNN COCO Cap 22.8 46.0 39.9 ViLD-ens [13] (ICLR ’22) R50-FPN CLIP ViT-B/32 Mask-RCNN - 27.6 59.5 51.3 RegionCLIP [159] (CVPR ’22) R50-C4 CLIP RN50 Faster-RCNN COCO Cap 26.8 54.8 47.5 RegionCLIP [159] (CVPR ’22) R50-C4 CLIP RN50 Faster-RCNN CC3M 31.4 57.1 50.4 Detic [160] (ECCV ’22) R50-FPN CLIP CenterNet2 IN-L 27.8 47.1 45.0 PB-OVD [190] (ECCV ’22) R50-FPN CLIP ViT-B/32 Mask-RCNN COCO Cap 30.8 46.1 42.1 HierKD [157] (CVPR ’22) R50-FPN CLIP ATSS COCO Cap 20.3 51.3 43.2 LocOV [192] (GCPR ’22) R50-FPN CLIP Faster-RCNN COCO Cap 28.6 51.3 45.7 OV-DETR [162] (ECCV ’22) R50-C4 CLIP Deformable DETR - 29.4 52.7 61.0 MEDet [262] (arXiv’22) R50-FPN CLIP Faster-RCNN COCO Cap 32.6 53.5 48.0 PromptDet [36] (ECCV ’22) R50-FPN CLIP Mask-RCNN LAION 31.4 29.1 29.7 VLDet [35] (ICLR ’23) R50-FPN CLIP CenterNet2 COCO Cap 32.0 50.6 45.8 F-VLM [165] (ICLR ’23) R50-FPN CLIP Mask-RCNN - 28.0 - 39.6 BARON [166] (CVPR ’23) R50-FPN CLIP Faster-RCNN - 34.0 60.4 53.5 BARON [166] (CVPR ’23) R50-FPN CLIP Faster-RCNN COCO Cap 42.7 54.9 51.7 OADP [182] (CVPR ’23) R50-C4 CLIP Faster-RCNN COCO Cap 30.0 53.3 47.2
@@
TABLE IV: Open vocabulary object detection results on LVIS 1.0 dataset.
@@
Method Vision Backbone Text Model Trainable Backbone Detector Extra Data $AP_{r}$ $AP_{c}$ $AP_{f}$ $AP_{all}$ ViLD-ens [13] (ICLR’22) R50-FPN CLIP ViT-B/32 Mask-RCNN - 16.6 24.6 30.3 25.5 RegionCLIP [159] (CVPR’22) R50-C4 CLIP RN50 Faster-RCNN CC3M 17.1 27.4 34.0 28.2 Detic [160] (ECCV’22) R50-FPN CLIP CenterNet2 IN-L 19.5 - - 30.9 OV-DETR [162] (ECCV’22) R50-C4 CLIP Deformable DETR - 17.4 25.0 32.5 26.6 MEDet [262] (arXiv’22) R50-FPN CLIP Faster-RCNN CC3M 22.4 - - 34.4 DetPro [161] (ECCV’20) R50-FPN CLIP ViT-B/32 Mask-RCNN - 19.8 25.6 28.9 25.9 PromptDet [36] (ECCV’22) R50-FPN CLIP Mask-RCNN LAION 21.4 23.3 29.3 25.3 VLDet [35] (ICLR’23) R50-FPN CLIP CenterNet2 CC3M 21.7 29.8 34.3 30.1 F-VLM [165] (ICLR’23) R50-FPN CLIP Mask-RCNN - 18.6 - - 24.2 BARON [166] (CVPR’23) R50-FPN CLIP Faster-RCNN - 22.6 27.6 29.8 27.6 OADP [182] (CVPR’23) R50-C4 CLIP Faster-RCNN - 21.7 26.3 29.0 26.6 GridCLIP [263] (arXiv’23) RN50 ViT-B/32 FCOS - 15.0 22.7 32.5 25.2 RegionCLIP [159] (CVPR ’22) R50x64-C4 CLIP RN50x64 Faster-RCNN CC3M 22.0 32.1 36.9 32.3 Detic [160] (ECCV’22) Swin-B CLIP CenterNet2 IN-L 23.9 40.2 42.8 38.4 VLDet [35] (ICLR’23) Swin-B CLIP CenterNet2 CC3M 26.3 39.4 41.9 38.1 F-VLM [165] (ICLR’23) R50x64-FPN CLIP Mask-RCNN - 32.8 - - 34.9
@@
4.1 Open Vocabulary Object Detection
@@
Settings. Open vocabulary object detection methods, such as OVR-CNN [12] and ViLD [13], have primarily focused on the COCO and LVIS datasets. The COCO dataset comprises 80 classes, with 48 considered base classes and 17 treated as novel classes. Any annotations not labeled with base classes are removed from the training data. Therefore, for open vocabulary object detection using the COCO dataset, there are 107,761 training images containing 665,387 bounding box annotations of base classes and 4,836 test images with 28,538 bounding box annotations of both base and novel categories. On the other hand, the LVIS dataset is designed for long-tail object detection tasks and includes a total of 1203 classes. There are 866 frequent and common ones that serve as base categories, while the remaining rare ones (377) act as novel categories.
@@
Results Comparison. In Tab. III & IV, we present the performance of current open vocabulary object detection methods on COCO and LVIS. BARON [166] outperforms other models on both datasets when using Faster-RCNN detector with the ResNet-50 backbone, achieving a high score of 42.7 $AP_{novel}$ on COCO and 22.6 $AP_{r}$ on LVIS. FVLM [165] achieves 28.0 $AP_{novel}$ on the COCO dataset and 18.6 $AP_{r}$ on LVIS using only frozen backbone and detection data. When utilizing the ResNet50x64 backbone, FVLM achieves a remarkable performance of 32.8 $AP_{r}$ on the LVIS dataset, surpassing other methods that employ the Swin-B backbone and additional data like CC3M and ImageNet. And for one-stage detector methods such as HierKD [157] and GridCLIP [263], there is still a significant gap compared to two-stage detectors.
TABLE V: Open vocabulary semantic segmentation performances under the self-evaluation setting. The metric is mIoU. “Harnomic” means the harmonic mean of the mIoU for base classes and the mIoU for novel classes. The best results are emphasized in bold.
@@
FreeSeg [216] (arXiv’23) RN-101 CLIP-ViT-B/16 42.2 49.1 45.3 91.8 82.6 86.9 - - -
@@
MaskCLIP [39] (arXiv’22) RN-50 CLIP-ViT-L/14 None 8.2 10.0 23.7 45.9 - -
@@
Han et al. [207] (arXiv’23) RN-101 CLIP-RN-50 None 3.5 7.1 18.8 45.2 83.2 -
@@
SAN [180] (CVPR’23) ViT-B/16 CLIP-ViT-B/16 None 10.1 12.6 27.5 53.8 - 94.0
SegCLIP [212] (arXiv’22) ViT-B/16 CLIP-ViT-B/16 CC3M - - - 24.7 - 52.6
PACL [211] (arXiv’22) ViT-B/16 CLIP-ViT-B/16 CC3M + CC12M + YFCC - - 31.4 50.1 - 72.3
X-Decoder [215] (arXiv’22) DaViT-B None FLD-900M 8.2 14.7 27.2 65.1 - 97.9
@@
4.2 Open Vocabulary Semantic Segmentation
Settings. There are two main settings in open vocabulary semantic segmentation, according to different train/test data choices. 1) Self-evaluation setting. Methods like ZegFormer [169] and MaskCLIP+ [170] train and test their models on the train/test split from the same dataset. For example, ZegFormer trains on base annotations on the COCO-Stuff training set and tests on the COCO-Stuff validation set. 2) Cross-evaluation setting. Methods like OpenSeg [37] and OVSeg [171] train their models on COCO-Stuff and test on other datasets, such as ADE20K, Pascal VOC, and Pascal Context. There is also a dataset split setting that splits a dataset into multiple folds and performs n-fold evaluations. We do not compare the results under this setting because only a few earlier works adopt it.
@@ 4.2 Open Vocabulary Semantic Segmentation
Evaluation Metrics. We adopt mean Intersection-over-Union (mIoU) to evaluate a model’s performance on open vocabulary semantic segmentation. Specifically, for the cross-evaluation setting, the IoU averaged on all the categories in the evaluation datasets is listed because all the classes can be seen as novel classes. For the self-evaluation setting, the testing dataset usually contains both base and novel classes. For a fair comparison, we only report mIoUs on novel classes, while highlighting that the number is in a different setting.
Results under the Self-evaluation Setting. Tab. V shows the open vocabulary semantic segmentation results under the self-evaluation setting. MaskCLIP+ [170] achieves the highest novel mIoU on COCO-Stuff, PASCAL-VOC, and PASCAL-Context, while in general, FreeSeg [216] has the best scores on base and harmonic mIoU. $*{}^{*}$ means including complicated crop-mask image preprocess (CLIP-Image encoder).
@@
TABLE VII: Open vocabulary instance segmentation performances on the COCO-Instances dataset. The best results are emphasized in bold.
@@
mask cls cap Base Novel Base Novel All
@@
CGG [38] (arXiv’23) RN-50 None None 46.8 29.5 46.0 28.4 41.4
TABLE VIII: Open vocabulary panoptic segmentation performances on the ADE20K dataset. No methods use extra data for training. The best results are emphasized in bold.
@@
MaskCLIP [39] (arXiv’22) RN-50 CLIP-ViT-L/14 15.1 13.5 18.3 70.5 19.2
@@
OPSNet [265] (arXiv’23) Swin-L CLIP-RN-50 17.7 15.6 21.9 54.9 21.6
TABLE IX: Open vocabulary panoptic segmentation performances on the COCO dataset. No methods use extra data for training. The best results are emphasized in bold.
@@
4.3 Open Vocabulary Instance Segmentation

### v4

@@ Appendix A Benchmark Results
This section systematically compares the different settings and methods in each task. For each setting, first, we introduce the datasets used in the evaluation. Next, we present the details of each setting separately. Then, we compare the results in detailed tables. We list the detailed results and settings for reference.
@@
TABLE IX: Open vocabulary object detection results on the COCO dataset. The evaluation metric is box mAP at IoU threshold 0.5 for novel and base classes, respectively. The best results are emphasized in bold.
@@
Method Vision Backbone Text Model Trainable Backbone Detector Extra Data $AP_{novel}$ $AP_{base}$ $AP_{all}$ OVR-CNN [11] (CVPR’21) RN50-C4 BERT Faster R-CNN COCO Cap 22.8 46.0 39.9 ViLD-ens [12] (ICLR ’22) RN50-FPN CLIP ViT-B/32 Mask R-CNN - 27.6 59.5 51.3 RegionCLIP [216] (CVPR ’22) RN50-C4 CLIP RN50 Faster R-CNN COCO Cap 26.8 54.8 47.5 RegionCLIP [216] (CVPR ’22) RN50-C4 CLIP RN50 Faster R-CNN CC3M 31.4 57.1 50.4 Detic [134] (ECCV ’22) RN50-FPN CLIP CenterNet2 IN-L 27.8 47.1 45.0 PB-OVD [119] (ECCV ’22) RN50-FPN CLIP ViT-B/32 Mask R-CNN COCO Cap 30.8 46.1 42.1 HierKD [224] (CVPR ’22) RN50-FPN CLIP ATSS COCO Cap 20.3 51.3 43.2 LocOV [238] (GCPR ’22) RN50-FPN CLIP Faster R-CNN COCO Cap 28.6 51.3 45.7 OV-DETR [218] (ECCV ’22) RN50-C4 CLIP Deformable DETR - 29.4 52.7 61.0 MEDet [336] (arXiv’22) RN50-FPN CLIP Faster R-CNN COCO Cap 32.6 53.5 48.0 PromptDet [45] (ECCV ’22) RN50-FPN CLIP Mask R-CNN LAION 31.4 29.1 29.7 VLDet [44] (ICLR ’23) RN50-FPN CLIP CenterNet2 COCO Cap 32.0 50.6 45.8 F-VLM [137] (ICLR ’23) RN50-FPN CLIP Mask R-CNN - 28.0 - 39.6 BARON [113] (CVPR ’23) RN50-FPN CLIP Faster R-CNN - 34.0 60.4 53.5 OADP [226] (CVPR ’23) RN50-C4 CLIP Faster R-CNN COCO Cap 30.0 53.3 47.2 CORA [219] (CVPR ’23) RN50 CLIP DAB-DETR - 35.1 35.5 35.4 EdaDet [337] (ICCV’23) RN50 CLIP-RN50 DETR - 37.8 57.7 52.5 CoDet [250] (NeurIPS’23) RN50 CLIP CenterNet2 COCO Caption 30.6 52.3 46.6 RO-ViT [251] (CVPR’23) ViT-B CLIP Mask R-CNN - 30.2 - 41.5 DITO [234] (arXiv’23) ViT-B/16 DITO Pretrain Faster R-CNN - 38.6 - 48.5 CLIPSelf [338] (ICLR’24) ViT-B/16 CLIP Mask R-CNN - 37.6 54.9 50.4 DST-Det [118] (arXiv’23) ViT-B/16 CLIP Mask R-CNN - 41.3 - - ProxyDet [339] (AAAI ’24) RN50-C4 CLIP ViT-B/32 Faster R-CNN - 30.4 52.6 46.8 LP-OVOD [223] (WACV ’24) RN50 CLIP Faster R-CNN - 40.5 60.5 55.2
@@
TABLE X: Open vocabulary object detection results on LVIS 1.0 dataset. The evaluation metric is mask mAP, and $AP_{r}$ represents the performance for novel categories.
@@
Method Vision Backbone Text Model Trainable Backbone Detector Extra Data $AP_{r}$ $AP_{c}$ $AP_{f}$ $AP_{all}$ ViLD-ens [12] (ICLR’22) RN50-FPN CLIP ViT-B/32 Mask R-CNN - 16.6 24.6 30.3 25.5 RegionCLIP [216] (CVPR’22) RN50-C4 CLIP RN50 Faster R-CNN CC3M 17.1 27.4 34.0 28.2 Detic [134] (ECCV’22) RN50-FPN CLIP CenterNet2 IN-L 19.5 - - 30.9 OV-DETR [218] (ECCV’22) RN50-C4 CLIP Deformable DETR - 17.4 25.0 32.5 26.6 MEDet [336] (arXiv’22) RN50-FPN CLIP Faster R-CNN CC3M 22.4 - - 34.4 DetPro [217] (ECCV’20) RN50-FPN CLIP ViT-B/32 Mask R-CNN - 19.8 25.6 28.9 25.9 PromptDet [45] (ECCV’22) RN50-FPN CLIP Mask R-CNN LAION 21.4 23.3 29.3 25.3 VLDet [44] (ICLR’23) RN50-FPN CLIP CenterNet2 CC3M 21.7 29.8 34.3 30.1 F-VLM [137] (ICLR’23) RN50-FPN CLIP Mask R-CNN - 18.6 - - 24.2 BARON [113] (CVPR’23) RN50-FPN CLIP Faster R-CNN - 22.6 27.6 29.8 27.6 OADP [226] (CVPR’23) RN50-C4 CLIP Faster R-CNN - 21.7 26.3 29.0 26.6 CORA [219] (CVPR’23) RN50 CLIP DAB-DETR - 22.2 - - EdaDet [337] (ICCV’23) RN50 CLIP RN50 DETR - 23.7 27.5 29.1 27.5 CoDet [250] (NeurIPS’23) RN50 CLIP CenterNet2 CC3m 23.4 30.0 34.6 30.7 ProxyDet [339] (AAAI ’24) RN50 CLIP ViT-B/32 Mask R-CNN IN-L 26.2 - - 32.5 LP-OVOD [223] (WACV ’24) RN50 CLIP Mask R-CNN - 19.3 29.4 26.1 26.2 RO-ViT [251] (CVPR’23) ViT-B CLIP Mask R-CNN - 28.0 - - 30.2 DITO [234] (arXiv’23) ViT-B/16 DITO Pretrain Mask R-CNN 32.5 - - - 34.0 CLIPSelf [338] (ICLR’24) ViT-B/16 CLIP Mask R-CNN - 25.3 21.8 29.1 25.2 DST-Det [118] (arXiv’23) ViT-B/16 CLIP Mask R-CNN - 26.2 - - - RegionCLIP [216] (CVPR ’22) RN50x64-C4 CLIP RN50x64 Faster R-CNN CC3M 22.0 32.1 36.9 32.3 Detic [134] (ECCV’22) Swin-B CLIP CenterNet2 IN-L 23.9 40.2 42.8 38.4 VLDet [44] (ICLR’23) Swin-B CLIP CenterNet2 CC3M 26.3 39.4 41.9 38.1 F-VLM [137] (ICLR’23) RN50x64-FPN CLIP Mask R-CNN - 32.8 - - 34.9 CLIPSelf [338] (ICLR’24) ViT-L/16 CLIP Mask R-CNN - 34.9 34.6 35.6 35.1 DST-Det [118] (arXiv’23) RN50x64-FPN CLIP Mask R-CNN - 34.5 - - -
@@
TABLE XI: Open vocabulary object detection results on V3Det [50] dataset. The metric is mAP on different IoU thresholds.
Method Vision Backbone Text Model Trainable Backbone Detector Extra Data $AP_{novel}$ $AP_{base}$ $AP_{all}$ RegionCLIP [216] (CVPR’22) RN50 CLIP Faster R-CNN - 3.1 22.1 12.6 Detic [134] (ECCV’22) RN50 CLIP CenterNet2 - 6.7 30.2 17.1 DST-Det [118] (arXiv’23) RN50 CLIP Faster R-CNN - 7.2 - - DST-Det [118] (arXiv’23) RN50x64 CLIP Faster R-CNN - 13.5 - -
A.1 Open Vocabulary Object Detection
@@
Settings. Open vocabulary object detection methods, such as OVR-CNN [11] and ViLD [12], have primarily focused on the COCO [5], LVIS [49] and V3Det [50] datasets. The COCO dataset comprises 80 classes, with 48 considered base classes and 17 treated as novel classes. Any annotations not labeled with base classes are removed from the training data. Therefore, for open vocabulary object detection using the COCO dataset, there are 107,761 training images containing 665,387 bounding box annotations of base classes and 4,836 test images with 28,538 bounding box annotations of both base and novel categories. On the other hand, the LVIS dataset is designed for long-tail object detection tasks and includes 1203 classes. 866 frequent and common ones serve as base categories, while the remaining rare ones (377) act as novel categories. The LVIS dataset is specifically designed for long-tail object detection tasks and consists of 1203 classes. Among these, there are 866 frequent and common categories that serve as base categories, while the remaining 377 rare ones serve as novel categories. The V3Det dataset is a vast vocabulary visual detection dataset that contains 13,204 categories, 243k images, and 1,753k bounding box annotations. In the open vocabulary setting, there are 6,709 base categories and 6,495 novel categories.
@@
Results Comparison. In Table IX, X, and XI, we present the performance of current open vocabulary object detection methods on COCO, LVIS, and V3Det. BARON [113] outperforms other models on both datasets when using Faster R-CNN detector with the ResNet-50 backbone, achieving a high score of 42.7 $AP_{novel}$ on COCO and 22.6 $AP_{r}$ on LVIS. FVLM [137] achieves 28.0 $AP_{novel}$ on the COCO dataset and 18.6 $AP_{r}$ on LVIS using only frozen backbone and detection data. When utilizing the ResNet50x64 backbone, F-VLM achieves a remarkable performance of 32.8 $AP_{r}$ on the LVIS dataset, surpassing other methods that employ the Swin-B backbone and additional data like CC3M and ImageNet. And for one-stage detector methods such as HierKD [224] and GridCLIP [340], there is still a significant gap compared to two-stage detectors.
TABLE XII: Open vocabulary semantic segmentation performances under the self-evaluation setting. The metric is mIoU. “Harmonic” means the harmonic mean of the mIoU for base classes and the mIoU for novel classes.
@@
FreeSeg [131] (CVPR’23) RN-101 CLIP-ViT-B/16 42.2 49.1 45.3 91.8 82.6 86.9 - - -
@@
MaskCLIP [48] (ICML’23) RN-50 CLIP-ViT-L/14 None 8.2 10.0 23.7 45.9 - -
@@
Han et al. [255] (ICCV’23) RN-101 CLIP-RN-50 None 3.5 7.1 18.8 45.2 83.2 -
@@
SegCLIP [126] (ICML’23) ViT-B/16 CLIP-ViT-B/16 CC3M - - - 24.7 - 52.6
PACL [261] (CVPR’23) ViT-B/16 CLIP-ViT-B/16 CC3M + CC12M + YFCC - - 31.4 50.1 - 72.3
PGSeg [342] (arXiv’23) ViT-S/16 None CC12M + RedCaps12M - - - 23.8 - 53.2
PnP-OVSS [343] (arXiv’23) ViT-L/14 BLIP-ViT-L None - - - 41.9 - 55.7
GEM [344] (arXiv’23) ViT-B/16 CLIP-ViT-B/16 None - - - 34.5 - 46.2
Self-SEG [345] (arXiv’23) Focal-T BLIP-ViT-L None 6.4 - - - 41.6 -
CaR [346] (arXiv’23) ViT-L/14 CLIP-ViT-B/16 None - - - 30.5 - 67.6
@@
SED [347] (arXiv’23) - CLIP-ConvNeXt-L None 13.9 22.6 35.2 60.6 - 96.1
SCAN [254] (arXiv’23) Swin-B CLIP-ViT-L/14 None 14.0 16.7 33.5 59.3 - 97.2
X-Decoder [130] (CVPR’23) DaViT-L None [348, 349, 350, 33] 9.2 16.1 29.6 64.0 - 97.7
OpenSeed [128] (ICCV’23) Swin-L None Objects365 [351] - - 23.4 - - -
HIPIE [352] (arXiv’23) ViT-H None Objects365 [351] - - 29.0 - - -
UOVN [353] (arXiv’23) RN-50 None [202, 354, 350, 355] 13.5 17.1 30.7 54.3 - -
A.2 Open Vocabulary Semantic Segmentation
Settings. There are two main settings in open vocabulary semantic segmentation, according to different train/test data choices. 1) Self-evaluation setting. Methods like ZegFormer [114] and MaskCLIP+ [120] train and test their models on the train/test split from the same dataset. For example, ZegFormer trains on base annotations on the COCO-Stuff training set and tests on the COCO-Stuff validation set with base and novel categories. 2) Cross-evaluation setting. Methods like OpenSeg [46] and OVSeg [121] train their models on COCO-Stuff and test on other datasets, such as ADE20K, Pascal VOC, and Pascal Context. There is also another setting that splits a dataset into multiple folds and performs n-fold evaluations. We do not compare the results under this setting considering only few works adopt it [14, 252].
@@ A.2 Open Vocabulary Semantic Segmentation
Evaluation Metrics. We adopt mean Intersection-over-Union (mIoU) to evaluate a model’s performance on open vocabulary semantic segmentation. Specifically, for the cross-evaluation setting, the IoU averaged on all the categories in the evaluation datasets is listed because all the classes can be seen as novel. For the self-evaluation setting, the testing dataset usually contains both base and novel classes. For a fair comparison, we only report mIoUs on novel classes while highlighting that the number is in a different setting.
Results under the Self-evaluation Setting. Tab. XII shows the open vocabulary semantic segmentation results under the self-evaluation setting. MaskCLIP+ [120] achieves the highest novel mIoU on COCO-Stuff, PASCAL-VOC, and PASCAL-Context, while in general, FreeSeg [131] has the best scores on base and harmonic mIoU. $*{}^{*}$ means including complicated crop-mask image preprocess (CLIP-Image encoder).
@@
TABLE XIV: Open vocabulary instance segmentation performances on the COCO-Instances dataset.
@@
mask cls cap $AP_{base}$ $AP_{novel}$ $AP_{base}$ $AP_{novel}$ $AP_{all}$
[…]
@@
CGG [47] (ICCV’23) RN-50 None None 46.8 29.5 46.0 28.4 41.4
TABLE XV: Open vocabulary panoptic segmentation performances on the ADE20K dataset. No methods use extra data for training.
@@
MaskCLIP [48] (ICML’23) RN-50 CLIP-ViT-L/14 15.1 13.5 18.3 70.5 19.2
@@
OPSNet [115] (ICCV’23) Swin-L CLIP-RN-50 17.7 15.6 21.9 54.9 21.6
TABLE XVI: Open vocabulary panoptic segmentation performances on the COCO dataset. No methods use extra data for training.
@@
A.3 Open Vocabulary Instance Segmentation